# https://zielon.github.io/gem/

### v1

@@ 4 Gaussian Eigen Model (GEM)
Figure 3: Samples of a GEM. We display samples for the first three components of the geometry eigenbasis of a GEM in the range of $[-3\sigma,3\sigma]$, showing diverse expressions. Note that GEM requires no parametric 3D face model like FLAME[37].
We build our statistical model for each modality of the Gaussians from $\mathbf{A}$ separately, specifically, we have individual bases for rotation, position, opacity, and scales. To compute the ensemble of eigenbasis we use Scikit-learn [51] and PCA [25] and define our model as an orthogonal coordinate system formed by the eigenvectors of the covariance matrices of each modality, formally we solve:
$(\mathbf{A}^{T}\mathbf{A})\mathbf{V}=\mathbf{V}\mathbf{\Lambda},$ (5)
where $\mathbf{A}^{T}\mathbf{A}\in\mathbb{R}^{M\times M}$ is the covariance matrix, $\mathbf{V}\in\mathbb{R}^{M\times M}$ is the matrix of eigenvectors and $\mathbf{\Lambda}\in\mathbb{R}^{M\times M}$ is the diagonal matrix of eigenvalues ($M=N\times 11\times H\times W$). Thus, per modality, we have:
$\mathbf{S}_{mod}=\mathbf{\bar{S}}_{mod}+\sum^{m}_{i=1}\lambda_{i}\mathbf{v}_{i},$ (6)
where $\lambda_{i}$ are coefficient ($\mathbf{\Lambda}_{:,i}$), $\mathbf{\bar{S}}_{mod}$ is a given modality mean, $\mathbf{v}_{i}$ are the spanning vectors of the corresponding orthonormal basis $\mathbf{V}_{:,i}$, and the number of principal components is set to $m=50$. 3 depicts samples of a GEM on the 3D position modality. For the remainder of the document, we omit the mod index for simplicity.
After estimating the initial basis $\mathbf{V}$ which might contain artifacts produced by the CNN network, we refine it using a photometric reconstruction loss on the training images for around 30k iterations employing the same objectives from Equation 4 to obtain $\mathbf{\hat{V}}$. To ensure that the basis stays orthonormal throughout this refinement, every 1k steps, we orthogonalize $\mathbf{\hat{V}}$ using QR decomposition such that $\mathbf{\hat{V}}^{T}\mathbf{\hat{V}}=\mathbf{\hat{V}}\mathbf{\hat{V}}^{T}=\mathbf{I}$. Example training PSNR errors w.r.t refinement steps 0k, 5k, and 30k are 34.75dB, 36.68dB, and 36.85dB, respectively, showing further improvement in the quality.
4.1 Image-based animation of GEM
Expressions for a GEM are fully defined by their coefficients. This is a similar idea to codec avatars introduced by Ma et al. [42], however, our approach does not need additional small regression MLPs. There are several ways to obtain the coefficients of a GEM, for example, one could employ analysis-by-synthesis-based optimization or regression. Analysis-by-synthesis is the backbone of current avatar methods, as they use photometric or depth-based face trackers to sequentially optimize the coefficients of the underlying 3DMMs like FLAME [67, 84, 17] which is typically slow. As a fast, but more imprecise alternative, regressors like DECA [9] or EMOCA [7] can be used which are built on a ResNet backbone and regress FLAME parameters directly from an image.
@@ 4.1 Image-based animation of GEM
As we aim for fast and expressive motion capturing, we build a personalized GEM regressor, that builds upon EMOCA, however, with several major modifications, see 4. First, we do not use FLAME parameters, instead, we use features extracted from a pre-trained EMOCA network denoted as $\Theta(\mathbf{I}_{i})$ where $\mathbf{I}_{i}$ is the current image. EMOCA’s architecture comprises two ResNet networks; one to extract expression features $\mathbf{f}_{expr}\in\mathbb{R}^{2048}$ and the second for shape $\mathbf{f}_{shape}\in\mathbb{R}^{2048}$, both are followed by final MLPs to regress corresponding FLAME parameters. In contrast, we remove the last hidden layer of the final MLP obtaining features $\in\mathbb{R}^{1024}$ and combine the regressed vectors into one $\mathbf{f}_{i}\in\mathbb{R}^{2\times 1024}$ vector. For training, we select five frontal cameras from NeRSemble [33]. Next, we build a PCA layer using the training frames similar to the one from our appearance model, however, with the difference that we use relative features $\mathbf{r}_{i}=\mathbf{f}_{i}-\mathbf{f}_{neutral}$ instead of meshes. The neutral reference frame $\mathbf{f}_{neutral}=\Theta(\mathbf{I}_{neutral})$ is selected manually from the video similar to Face2Face [67]. During training, for each frame, we project $\mathbf{r}_{i}$ onto the PCA manifold to restrict and regularize training using the first 50 components. The projected coefficients are passed through a small MLP that produces a vector of GEM coefficients per modality, denoted as $\boldsymbol{p}_{i}\in\mathbb{R}^{M}$:
@@
$\begin{split}\mathbf{r}_{i}&=\Theta(\mathbf{I}_{i})-\mathbf{f}_{neutral},\\ \boldsymbol{p}_{i}&=3\sigma\cdot\tanh(\textrm{MLP}(\Gamma_{\text{PCA}}(\mathbf{r}_{i}))),\end{split}$ (7)
where $\Gamma_{\text{PCA}}$ is the relative PCA model and $M$ is the selected amount of linear basis. The MLP has three hidden layers with 256 neurons each and ReLU activations. We use a scaled $\tanh$ activation function for the output, to restrict the prediction to be in $[-3\sigma,3\sigma]$. The final 3D model is obtained by $\mathbf{\bar{S}}+\boldsymbol{p}_{i}^{T}\mathbf{\hat{S}}$ which is splatted using 3DGS. We follow the objective function from Equation 4 for supervision on the color reproduction with an additional L1 regularization w.r.t. the corresponding coefficients. We use AdamW with a learning rate $1\mathrm{e}{-3}$ and train it for 200k iterations with batch size one on a single Nvidia Quadro A6000 which takes around 4h.
@@ 5 Results
Figure 5: Novel view synthesis. Both, the network and GEM, show better performance on novel views, especially, in the region of the mouth interior and wrinkles. In this experiment, we are following the evaluation of Gaussian Avatars [53] and demonstrate novel viewpoint generation. GEM is obtained throughout analysis-by-synthesis fitting [3, 67]. Note that the expressions are seen during training.
We evaluate our appearance model against state-of-the-art methods on the NeRSemble [33] dataset. We use the tracked meshes provided by Qian et al. [53] and the synchronized images from 16 cameras with a resolution of $802\times 550$. Our main baselines are Gaussian Avatars (GA[53] which is neural network-free (Gaussians are attached to the FLAME model), our implementation of Animatable Gaussians (AG)  [38] which is based on CNN-predicting Gaussian maps, and INSTA [84] which uses dynamic NeRF [43]. For the comparison, we present both of our appearance models, the StyleUNet-based architecture (Ours Net) and the distilled linear eigen model (Ours GEM) which we evaluate using analysis-by-synthesis fitting to the target images following [85, 67, 3]. Additionally, we show cross-reenactment results based on relative FLAME parameters transfer [84] as well as our coefficients regressor versus other baselines which are using FLAME meshes regressed by DECA [9] instead of the ground truth ones from [53]. All of the methods are evaluated using several image space metrics on novel expressions and novel views, following the test and novel-view split from Qian et al. [53]. For all of our experiments with eigenbasis, we use 50 components distilled from $256^{2}$ textures which give around 60k active Gaussians. Animatable Gaussians [38] use a similar amount of primitives for front and back textures and Gaussian Avatars [53] typically use around 100k Gaussians.
Table 1: Evaluation on novel expressions and views shows improved results of our GEM optimized using analysis-by-synthesis compared to others. See 6 for the corresponding qualitative results.
@@
Ours GEM 32.6781 0.0675 0.9633 0.0069
Ours Net 29.2454 0.0777 0.9448 0.0096
@@
Table 2: Novel viewpoint evaluation on a withhold camera, corresponding to Figure 5. Note that the expressions have been seen during training, and only the view is new, the same as in GA [53].
Method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ L1 $\downarrow$
Ours GEM 33.5528 0.0678 0.9662 0.0061
Ours Net 32.4622 0.0713 0.9617 0.0067
AG [38] 32.4166 0.0712 0.9614 0.0066
GA [53] 31.3197 0.0786 0.9567 0.0075
INSTA [84] 27.7786 0.1232 0.9294 0.0163
@@ 5.1 Image Quality Evaluation
To evaluate our method, we measure the color error in the image space using the following metrics: PSNR (dB), LPIPS [77], L1 loss, and structural similarity (SSIM). We follow the evaluation scheme from Gaussian Avatars [53], using their train and validation split. The evaluation of GEM was generated by sequentially fitting the coefficients to each image using the same objectives from Equation 4. Note that the baselines use the FLAME model with offsets for the tracking, while GEM can directly be used for modeling.
Table 2 presents results on novel expressions evaluated on all 16 cameras. Both the quantitative and qualitative results depicted in Figure 6 show that our PCA model produces fewer artifacts, especially for regions like teeth or facial wrinkles. Table 2 contains an evaluation where we measure errors on novel viewpoints. The results show that our CNN-based appearance model outperforms other neural methods, and, our linear eigenbasis GEM achieves the highest quality due to the ’direct’ analysis-by-synthesis which can fully utilize the expressiveness and grade of our photo-realistic appearance model and is not restricted by 3DMMs like FLAME. Moreover, Figure 5 shows qualitative results of our method on novel views. As can be seen, we better capture high-frequency details, pose-dependent wrinkles, and self-shadows - something which is not possible for methods like Gaussian Avatars [53] or INSTA [84] since they either do not use pose-dependent neural networks or limit the conditioning to a small region only.
Refer to captionGround TruthOurs GEMOurs NetGA [53]AG [38]INSTA [84]
Figure 6: Novel view and expression synthesis. Our Gaussian Eigen Model shows better results in regions like teeth, wrinkles, and self-shadows compared to other methods that struggle with artifacts.
5.2 Cross-person Reenactment Evaluation
Facial cross-person reenactment transfers expressions from the source actor to the target actor. Most of the methods require tracked meshes obtained by fitting the FLAME model for each frame of the source actor sequence. From this tracking, the expression parameters are extracted and the relative change is applied to the neutral target face [67], see Figure 7. While this is an effective approach, it requires having the tracked mesh in a consistent mesh topology, which is a computationally intensive problem, especially in multi-view capture setups.
Table 3: Cross-re-enactment evaluation employing EmoNet features and FID score. *FPS omits the EMOCA forward pass time, which adds approximately 21 FPS to the inference.
Method $\textbf{E}_{feat}$cos $\uparrow$ $\textbf{E}_{feat}$ $\mathcal{L}_{1}$downarrow FID $\downarrow$ *FPS $\uparrow$
@@ 5.2 Cross-person Reenactment Evaluation
An alternative way is to use a (monocular) regressor like EMOCA [7] which, in real-time, can predict FLAME parameters for a single frame. We showcase this in comparison to our GEM which is driven by our image-based regressor, see Figure 8. As can be seen in both experiments, our network-based and GEM can produce sharp results. The baseline methods have problems extrapolating to the new expressions, showing severe artifacts in the appearance. Our personalized approach effectively regularizes the regressed coefficients such that the predicted avatar stays in the training distribution effectively avoiding artifacts like in INSTA or Gaussian Avatars.
@@
Refer to captionGround TruthOurs NetGA. [53]AG. [38]INSTA [84]
Refer to captionGround TruthOurs GEMOurs NetGA [53]AG [38]INSTA [84]
@@
5.3 GEM Compression Ablation
Refer to captionGround Truth#10#30#50
Figure 9: Compression error depending on the number of used eigenbasis components. The heatmaps show the photometric $\ell_{1}$-error for 10, 30, and 50 components using $128^{2}$ Gaussian maps. Full evaluation in the supp mat.
In this section, we perform ablation of our linear GEM model in several aspects. First, we measure the memory size of such a model depending on the number of selected components compared to the size of the regressed texture. Secondly, we are interested in the compression error introduced by the projection on different amounts of principal components. Table 4 shows how the size in MB changes depending on the amount of basis and the source texture size. For instance, our smallest model weighs as little as 7MB using only 10 components of the eigenbasis. This is almost 12 times less than our smallest CNN-based model and almost 70 times less than Animatable Gaussians [38]. In contrast to neural networks, we can easily trade quality over size which is very useful in the context of different commodity devices with reduced compute capabilities. Table 4 presents how compression affects the quality of reconstruction. We evaluated $\sim 1$k frames for a single actor under a novel view. As expected using only 10 first components impacts the quality the most, however, the results are still high-quality as can be seen in Figure 9. As we previously discussed, to capture wrinkles and high-frequence details one needs to employ neural networks. Gaussian Avatars [53] offers a small size of the stored Gaussians cloud, ranging from 5MB, and 14MB without the FLAME model for $128^{2}$, $256^{2}$ Gaussians, respectively. However, the quality of reconstruction lacks wrinkle details and sharpness as can be seen in Figure 10. Note that our model does not require a FLAME model that weights an additional 90MB compared to Gaussian Avatars [53]; Table 4 presents our absolute size values.

### v2

@@ 3.1 Gaussian Eigen Model (GEM)
Figure 3: Samples of a GEM. We display samples for the first three components of the position $\mathbf{k}_{\phi}$ eigenbasis of a GEM, showing diverse expressions. Note that GEM requires no parametric 3D face model like FLAME[40].
@@
Refer to captionGround TruthOurs GEMOurs NetGA [57]AG [41]INSTA [88]
Figure 5: Novel view synthesis. Both, our CNN and GEM show better performance on novel views, especially, in the region of the mouth interior and wrinkles. In this experiment, we are following the evaluation of Gaussian Avatars [57] and demonstrate novel viewpoint generation. GEM is obtained throughout analysis-by-synthesis fitting [3, 71]. Note that the expressions are seen during training.
3.2 Image-based Animation
Expressions for a GEM are fully defined by their coefficients $\mathbf{k}_{\theta}$, $\mathbf{k}_{\phi}$, $\mathbf{k}_{\theta}$ and $\mathbf{k}_{\sigma}$. This is a similar idea to codec avatars [45], however, our approach does not need additional pixel shaders in the form of a small regression MLP. There are several ways to obtain the coefficients of a GEM, for example, one can employ analysis-by-synthesis-based optimization or regression. Analysis-by-synthesis [3] is the backbone of current avatar methods, as they use photometric or depth-based face trackers to sequentially optimize the coefficients of the underlying 3DMMs like FLAME [71, 88, 19] which is typically slow. As a fast, but more imprecise alternative, regressors like DECA [10] or EMOCA [7] can be used which are built on a ResNet backbone and regress FLAME parameters directly from an image.
@@ 3.2 Image-based Animation
As we aim for fast and expressive motion capturing, we build a personalized GEM regressor, that utilizes a pretrained EMOCA model. We apply several modifications to the EMOCA model, see 4. We use intermediate features of the pre-trained EMOCA network denoted as $\Theta(\mathbf{I}_{i})$ where $\mathbf{I}_{i}$ is the current image. EMOCA’s architecture comprises two ResNet networks; one to extract expression features $\mathbf{f}_{expr}\in\mathbb{R}^{2048}$ and the second for shape $\mathbf{f}_{shape}\in\mathbb{R}^{2048}$, both are followed by final MLPs to regress corresponding FLAME parameters. As we do not rely on FLAME, we remove the last hidden layer of the final MLP obtaining two feature vectors which we combine into one $\mathbf{f}\in\mathbb{R}^{2\times 1024}$ vector. For these features, we build a PCA layer with a basis denoted as $\mathbf{\hat{R}}$ using the training frames from five frontal cameras of NeRSemble. Note that we use relative features $\mathbf{r}=\mathbf{f}-\mathbf{f}_{neutral}$ in this PCA layer. The neutral reference frame $\mathbf{f}_{neutral}=\Theta(\mathbf{I}_{neutral})$ to compute these relative features is selected manually from the video, similar to Face2Face [71]. During training, for each frame, we project $\mathbf{r}$ onto the PCA manifold using the first $50$ principal components to restrict and regularize training. Finally, we use their corresponding PCA coefficients:
$\begin{split}\mathbf{\kappa}&=(\mathbf{r}-\mathbf{\bar{R}})\mathbf{\hat{R}}^{T},\\ \end{split}$ (3)
where $\mathbf{\bar{R}}$ is the relative PCA model mean. The projected coefficients are passed through a small MLP that produces a vector of GEM coefficients $\mathbf{k}=\{\mathbf{k}_{\theta},\mathbf{k}_{\phi},\mathbf{k}_{\theta},\mathbf{k}_{\sigma}\}$:
@@
$\begin{split}\mathbf{k}&=3\cdot\sigma_{k}\cdot\tanh(\textrm{MLP}(\mathbf{\kappa})).\end{split}$ (4)
The MLP has three hidden layers with 256 neurons each and ReLU activations. We use a scaled $\tanh$ activation function for the output to restrict the prediction to be in $[-3\cdot\sigma_{k},3\cdot\sigma_{k}]$, $\sigma_{k}$ being the respective standard deviation of the coefficients $\mathbf{k}$, obtained from the PCA. The final primitives are obtained by Eq. 1 and splatted using 3DGS.
@@ 4 Results
Figure 6: Novel view and expression synthesis. Our Gaussian Eigen Models for Human Heads shows better results in regions like teeth, wrinkles, and self-shadows compared to other methods that struggle with artifacts.
We evaluate GEM on the NeRSemble [33], where tracked meshes [57] and synchronized images from 16 cameras with a resolution of $802\times 550$ are available. Our baselines are Gaussian Avatars (GA[57] which is neural network-free (Gaussians are attached to the FLAME model), our implementation of Animatable Gaussians (AG)  [41] which is based on CNN-predicting Gaussian maps, and INSTA [88] which uses dynamic NeRF [46]. For the comparison, we present both of our appearance models, the StyleUNet-based architecture (Ours Net) and the distilled linear Eigen model (Ours GEM) which we evaluate using analysis-by-synthesis fitting to the target images following [89, 71, 3]. Additionally, we present cross-reenactment results based on our coefficient regressor, compared to the baselines that use FLAME meshes regressed by DECA [10]. Relative expression transfer based on ground truth meshes [57] can be found in the supp. mat. All of the methods are evaluated using several image space metrics on novel expressions and novel views, following the test and novel-view split of Qian et al. [57]. For our GEM models, we use 50 components distilled from $256^{2}$ textures which give around 60k active Gaussians. Animatable Gaussians [41] uses a similar amount of primitives for front and back textures and Gaussian Avatars [57] around 100k Gaussians.
@@
AG [41] 32.4166 0.0712 0.9614 0.0066
GA [57] 31.3197 0.0786 0.9567 0.0075
INSTA [88] 27.7786 0.1232 0.9294 0.0163
Ours Net 32.4622 0.0713 0.9617 0.0067
Ours GEM 33.5528 0.0678 0.9662 0.0061
Table 1: Novel viewpoint evaluation is conducted on a withhold camera from the 16 cameras used for training. Note that the expression has been seen during training, and only the view is new.
Method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ L1 $\downarrow$
@@
Ours Net 29.2454 0.0777 0.9448 0.0096
Ours GEM 32.6781 0.0675 0.9633 0.0069
Table 2: Evaluation on novel expressions and views show improved results of GEM optimized using analysis-by-synthesis compared to others. Figure 6 shows the corr. qualitative results.
Refer to captionSource ActorOurs GEMOurs NetGA [57]AG [41]INSTA [88]
@@ 4.1 Image Quality Evaluation
To evaluate our method, we measure the color error in the image space using the following metrics: PSNR (dB), LPIPS [82], L1 loss, and structural similarity (SSIM). We follow the evaluation scheme from Gaussian Avatars [57], using their train and validation split. The evaluation of GEM was generated by sequentially fitting the coefficients to each image using photometric objectives. Note that the baselines use the FLAME model with offsets for the tracking, while GEM can directly be used for tracking.
Table 2 presents results on novel expressions evaluated on all 16 cameras. Both the quantitative and qualitative results depicted in Figure 6 show that our PCA model produces fewer artifacts, especially for regions like teeth or facial wrinkles. Table 1 contains an evaluation where we measure errors on novel viewpoints. The results demonstrate that our CNN-based appearance model outperforms other neural methods, while our linear eigenbasis GEM achieves the highest quality. This is due to the ’direct’ analysis-by-synthesis approach, which fully leverages the expressiveness and detail of our photorealistic appearance model, without the limitations imposed by 3DMMs such as FLAME. Moreover, Figure 5 shows qualitative results of our method on novel views. As can be seen, we better capture high-frequency details, pose-dependent wrinkles, and self-shadows - something which is not possible for methods like Gaussian Avatars [57] or INSTA [88], since they either do not use expression-dependent neural networks or limit the conditioning to a small region only.
Method $\textbf{E}_{feat}$cos $\uparrow$ $\textbf{E}_{feat}$ $\mathcal{L}_{1}$downarrow FID $\downarrow$ FPS $\uparrow$
@@
Table 3: Cross-reenactment evaluation employing EmoNet features and FID score.
4.2 Cross-person Reenactment Evaluation
Facial cross-person reenactment transfers expressions from the source actor to the target actor. For this, the baseline methods require tracked meshes obtained by fitting the 3DMM model for each frame of the source actor sequence. As an alternative to optimization-based tracking, a (monocular) regressor like EMOCA [7], can predict such tracked meshes in real-time. We demonstrate this in Figure 7, where GEM is driven by our image-based regressor and the others by EMOCA. As shown, our network-based method and GEM produce sharp results, while the baseline methods struggle to extrapolate to new expressions, displaying severe artifacts in appearance. Our approach effectively regularizes the regressed coefficients, ensuring that the predicted avatar remains in the training distribution and thereby avoids artifacts seen in INSTA or Gaussian Avatars.
@@ 4.2 Cross-person Reenactment Evaluation
Refer to captionGround Truth#10#30#50
Figure 8: Compression error depending on the number of used principle components in GEM. The heatmaps show the photometric $\ell_{1}$-error for 10, 30, and 50 components using $128^{2}$ Gaussian maps. See suppl. doc. for additional evaluations.
4.3 GEM Ablation Studies
We are interested in the compression error introduced by the projection on different amounts of principal components used in GEM, also concerning the memory consumption. For instance, our smallest model weighs as little as 7MB using only 10 components of the eigenbasis. This is almost 12 times less than our smallest CNN-based model and almost 70 times less than Animatable Gaussians [41]. In contrast to neural networks, we can easily trade quality over size which is very useful in the context of different commodity devices with reduced compute capabilities. Table 4 presents how compression affects the quality of reconstruction, where we evaluate a sequence with $\sim 1$k frames for a single actor under a novel view. As expected, using only 10 components impacts the quality the most, however, the results are still of high quality, see Figure 8.
Gaussian Avatars [57] offers a small size of the stored Gaussians cloud, ranging from 5MB, and 14MB without the FLAME model for $128^{2}$, $256^{2}$ Gaussians, respectively. However, the quality of reconstruction lacks wrinkle details and sharpness as can be seen in Figure 12. In comparison to Gaussian Avatars [57], our model does not require FLAME during inference which is an additional 90MB.
Refer to captionGround TruthOurs #30 $128^{2}$ [57]$128^{2}$ [57]$256^{2}$
Figure 9: The quality comparison to Gaussian Avatars. Note that we do not need an additional FLAME model which weighs 90MB.
@@ 4.3 GEM Ablation Studies
5 Discussion
We design a universal method capable of distilling 3DGS-based avatar solutions into a lightweight representation, GEM, provided that normalized input across training frames is available. The only requirement to successfully distill GEM is to have a dataset with Gaussian-image pairs across the training sequences. Our results show that a compact representation of the linear basis produces state-of-the-art results in terms of quality and speed. Note that to achieve wrinkle-level details, the generator itself has to produce high-quality outputs. Our distillation technique can be applied to existing methods, making them lightweight and compact. GEM is well-suited for commodity devices, generating Gaussian primitives by a simple linear combination of the basis vectors. This potential has promising implications for tasks like holoportation, audio-driven avatars, and virtual reality.
The PCA-based GEM models have a global extent which is useful for some applications, but it also means that we cannot control local changes and produce more combinations of local features. Thus, further work could include incorporating a localized PCA basis [48] for better avatar control, which could potentially enable a wider range of expressions outside the training set. Another limitation is the personalization, for new subjects a new representation has to be learned from multi-view data. An interesting future avenue is to create a statistical model across subjects.
6 Conclusion
We have proposed Gaussian Eigen Models for Human Heads, a linear appearance model that represents photo-realistic head avatars. The simplicity of this appearance model results in massively reduced compute requirements in comparison to CNN-based avatar methods. Moreover, the lightweight representation could improve the management, sharing, and applicability of avatars. Our distillation approach can be applied to existing methods, making them available for compression. We demonstrate how GEMs can be used in scenarios like self-reenactment and cross-person animation, even in real-time.
Acknowledgement
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting WZ. JT is supported by the Google Gift Award.
Appendix A 3D Gaussian Splatting Preliminaries
Appendix B Appearance Maps Generator
Similarly to StyleAvatar [74] and Animatable Gaussians [41], we use a StyleGAN-based [28] encoder and decoder network for this image translation. However, in contrast to Animatable Gaussians [41], we propose a more lightweight image translation pipeline, where we reduce the number of encoders from three to two and the number of decoders from six to two, and decrease the size of the StyleGAN [28] decoder. To efficiently use the 2D map space, we do not use projective textures from meshes [74], but a UV parametrization which reduces half of the decoders in comparison to StyleAvatar. Moreover, we do not use triplets of the encoder-decoder, as we combine all properties of the Gaussians into one map $\boldsymbol{G}_{i}$ following ASH [51]. Therefore, we define our network as follows:
The final Gaussian map $\boldsymbol{G}_{i}\in\mathbb{R}^{11\times H\times W}$ is obtained by applying the deltas to the canonical Gaussians $\boldsymbol{G}$ [41, 90]. The deformed position of 3D means is computed as $\boldsymbol{T}_{i}(\boldsymbol{M}+\boldsymbol{D}_{i_{0:3}})$. One major distinction compared to AG [41] is the way how the transformation from the canonical space to the deformed space is handled. We employ deformation gradients, following Sumner et al. [64]. This approach allows for greater flexibility regarding input meshes, provided they maintain full correspondence.
Given a mesh $\boldsymbol{M}_{PCA_{i}}$ for the frame $i$, we define the deformation gradients as $\mathbf{J}_{j}=\hat{\mathbf{E}}_{j}\mathbf{E}_{j}^{-1}$, where $\hat{\mathbf{E}}_{j}\in\mathbb{R}^{3\times 3}$ and $\mathbf{E}_{j}\in\mathbb{R}^{3\times 3}$ contain the Frenet frame (tangent, bi-tangent, normal) of the triangle $j$ defined in deformed and canonical spaces, respectively. Using these deformation gradients and the known correspondences between the Gaussian map and the meshes, we transform the Gaussians from the canonical space to the deformed space.
Note that our color map C is static and does not model view-dependent effects; this means that we force the network to recover globally consistent colors for each Gaussian similar to a texture in the classic 3DMM. Therefore, $\boldsymbol{G}_{i}$ must model the wrinkles and self-shadows.
B.1 Image-based Coefficents Regressor
Table 5 shows an ablation study in the context of input to our MLP regressor which predicts GEM coefficients. Figure 11 provides an additional qualitative comparison. Using a pre-trained decoder such as EMOCA demonstrates strong potential for cross-reenactment by leveraging robust priors. Future work will explore further methods for image-based control of GEM, with a potential approach being the incorporation of additional modalities, such as sound, into the EMOCA-based regressor.
Ablation L1 $\downarrow$ LIPIS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Ours 0.0181 0.1171 24.2997 0.9134
Absolute 0.0181 0.1172 24.3065 0.9130
FLAME 0.0181 0.1169 24.2910 0.9130
EMOCA [7] 0.0180 0.1165 24.2715 0.9132
DECA [10] 0.0184 0.1189 24.2181 0.9127
Table 5: We performed an ablation study of our regressor using self-reenactment tasks. In this study, we tested configurations that used only EMOCA or DECA features, as well as a version where the EMOCA regressed FLAME expressions. Lastly, we show the effect of using absolute features instead of the relative ones used in Ours (features relative to the neutral face).
Refer to captionSourceRelativeDECA [10]AbsoluteEMOCA [7]FLAME [40]
Figure 11: Our experiments show that using pre-trained regressor like EMOCA [7] and DECA [10] work well for driving our GEM model. In this context, DECA and EMOCA refer to using either of the regressed feature vectors. FLAME represents the expression vectors regressed by EMOCA. Relative and absolute denote whether the EMOCA + DECA features are used directly or as relative changes from a neutral face.
B.2 CNN Training Details
where $\omega=0.2$, $\zeta=0.0075$ (after 150k iterations steps and zero otherwise), $\mathcal{L}_{\text{D-SSIM}}$ is a structural dissimilarity loss, and $\mathcal{L}_{\text{VGG}}$ is the perceptual VGG loss. $\mathcal{L}_{Reg}$ regularizes position offsets $\Delta_{pos_{j}}$ and scales $s_{scale_{j}}$ to stay small w.r.t. the input mesh. We train our model for $10^{6}$ steps using Adam [30] with a learning rate $5\mathrm{e}{-4}$ and a batch size of one which takes around 10h on a Nvidia RTX4090.